OpenStudy (anonymous):

for f(x)=ln(e^(3x) + e^(-3x)) find f'(x) and evaluate at x=ln2

OpenStudy (sidsiddhartha):

simply use chain rule

OpenStudy (sidsiddhartha):

u know how to use it?

OpenStudy (anonymous):

i'm not to sure can you guide me through?

OpenStudy (sidsiddhartha):

take a basic example\[A=[f(g(x))]\] now diffrentiating it with respect to x by chain rule--\[\frac{ dA }{ dx }=f'(g(x))*g'(x)\]

OpenStudy (anonymous):

would it be \[\frac{ dA }{ dx }=\frac{ -9e ^{-9x} }{ e ^{3x}+e ^{-3x} }\]

OpenStudy (sidsiddhartha):

no fisrt do a basic problem \[A=\ln(x^2)\] find the derivative

OpenStudy (anonymous):

\[\frac{ 1 }{ x^2 }\]

OpenStudy (sidsiddhartha):

yes here comes the problem ,here u have to use chain rule \[\frac{ dA }{ dx }=\frac{ d }{ dx } \ln(x^2)=\frac{ d }{ dx }\ln(x^2)*\frac{ d }{ dx }(x^2)=\frac{ 1 }{ x^2 }*2x\]

OpenStudy (sidsiddhartha):

getting this?

OpenStudy (anonymous):

yep

OpenStudy (jhannybean):

* treat \( (e^{3x} + e^{-3x})\) as the \(x^2\) you just learned how to take the derivative of from @sidsiddhartha :)

OpenStudy (sidsiddhartha):

yup can u do it now?

OpenStudy (anonymous):

its like this right? \[\frac{ 1 }{ e ^{-9x} }\times-9e ^{-9x}\]

OpenStudy (jhannybean):

When simplified, what does this give you?

OpenStudy (jhannybean):

\[\ \frac{d}{dx} (e^{3x}) = (e^{3x}) \cdot \frac{d}{dx} (3x)\] that is what I meant, sorry.

OpenStudy (anonymous):

\[3e ^{3x}\]

OpenStudy (jhannybean):

Good.

OpenStudy (jhannybean):

And \[\ \frac{d}{dx} (e^{-3x}) = e^{-3x} \cdot \frac{d}{dx} (-3x)\]

OpenStudy (jhannybean):

What about this one? :)

OpenStudy (anonymous):

\[-3e ^{-3x}\]

OpenStudy (jhannybean):

Awesome.

OpenStudy (anonymous):

so answer to first part is \[\frac{ 3e ^{3x}-3e ^{-3x} }{ e ^{3x}+e ^{-3x} }\]

OpenStudy (jhannybean):

You're very very close.

OpenStudy (jhannybean):

I'll write up my understanding of it and you can try following my logic. Ok? :)

OpenStudy (anonymous):

alright

OpenStudy (jhannybean):

So now that you've found the derivatives of both of those, we can find the derivative of the entire function. \[ \frac{d}{dx} \ln(e^{-3x} +e^{3x})\] By treating \(e^{-3x} + e^{3x}\) as \(x\) , we can think of this derivative as \(\frac{d}{dx} \ln(x)\) as \(\frac{1}{x}\). We have found both of our derivatives for the inner function, which is good, but I think there is an easier way to go about solving this instead of finding the derivatives separately and then putting it all together.\[\ \frac{d}{dx} \ln(e^{-3x} +e^{3x})= \frac{1}{(e^{-3x} +e^{3x})} \cdot \frac{d}{dx}(e^{-3x} +e^{3x})\]Taking the derivative of the inside can be a little tricky, but what if we made the power's of e positive instead?

OpenStudy (jhannybean):

We would have \[\ \frac{1}{(e^{-3x}+e^{3x})}\cdot \left[-\frac{3}{e^{3x}}+3e^{3x}\right] \]What I did was I made \(e^{-3x}\) into a +ve fraction by setting it over 1, and then using the chain rule to take the derivative of the inner function, like what we did before, and i simply took the derivative of the other function as you also had done earlier.

OpenStudy (jhannybean):

Now focusing solely on the function in brackets, we need to write it so that we have a common denominator to combine the function altogether. we can multiply the top and bottom of \(3e^{3x}\) by \(e^{3x}\).\[\left[ -\frac{3}{e^{3x}} + \frac{3e^{3x}\cdot e^{3x}}{e^{3x}}\right] = \frac{-3 +3e^{3x+3x}}{e^{3x}}=\frac{-3+3e^{6x}}{e^{3x}}=\frac{3(-1+e^{6x})}{e^{3x}}\]

OpenStudy (jhannybean):

Following so far? :\

OpenStudy (anonymous):

yep. i know it can be a very big hassle typing it out thanks for everything so far =)

OpenStudy (jhannybean):

Alright :) This solves the DERIVATIVE part of the function. Now we can concentrate on simplifying the major function as well. \[\frac{1}{(e^{-3x} +e^{3x})}\]rewriting this function, we can use the same method we used in finding the derivative. \[\large \frac{1}{\frac{1}{e^{3x}}+e^{3x}}=\frac{1}{\frac{1+e^{6x}}{e^{3x}}}\]And now we... combine EVERYTHING together. Oh god.

OpenStudy (jhannybean):

\[\large \frac{\frac{3(-1+e^{6x})}{e^{3x}(1+e^{6x})}}{e^{3x}}\]And cancelling the \(e^{3x}\) we get \[\frac{3(-1+e^{6x})}{(1+e^{6x})}\]

OpenStudy (jhannybean):

And that's how you solve this function :))!

OpenStudy (anonymous):

oh god if that is one question on a previous midterm we had to do then i am so worried about my midterm. we have 5 pages of questions and that was only 1 question.

OpenStudy (jhannybean):

Oh, don't stress. Just take it one step at a time and review each step as you do it.

OpenStudy (anonymous):

we have 60 mins for it =( a midterm i did before i ran out of time

OpenStudy (jhannybean):

Yeah... I feel that way when I take my calculus exams. you get used to working out hard problems really fast.

OpenStudy (anonymous):

i feel so unprepared

OpenStudy (jhannybean):

Oh, don't feel that way! I suggest retrying this problem over again and seeing if you can solve it on your own.

OpenStudy (jhannybean):

Good luck! :)

OpenStudy (anonymous):

thanks